Write an equation for a line perpendicular to y = -2x-5 and passing through the point (-6,-4)Step by step please.Thanks.

Answer by Cromlix(4381)

First obtain the slope of the
line y = -2x -5.
This = -2
Now lines that are perpendicular to one
another have slopes that multiply together
to give -1
So, slope 1 * slope 2 = -1
We have -2, so the other line is 1/2
as -2 * 1/2 = -1
We refer to the slope as the gradient
or as the letter 'm'
To set up the equation of the new line
we use the equation;
y - b = m(x - a)
We will be using coordinates (-6,-4)
a = -6, b = -4 and m = 1/2
y -(-4) = 1/2 (x -(-6)
y + 4 = 1/2(x + 6) (Multiply out)
y + 4 = 1/2x + 3
y = 1/2x + 3 - 4 (Bring 4 over (change sign to -ve))
y = 1/2x - 1
You can multiply thro by 2 to remove fractions.
2y = x - 2.
It just depends on how your teacher wants it displayed.
Hope this helps
:-)